# Discrete Linear Canonical Transform on Graphs: Uncertainty Principle and Sampling

### v2

@@ 1 Introduction
Advancements in information and communication technologies have led to the integration of various types of data in fields such as social networks, recommendation systems, medical image processing, and bioinformatics. Unlike traditional time series or digital images, these datasets often exhibit complex and irregular structures, posing challenges for conventional signal processing tools [1, 2]. Over the past decade, significant progress has been made in the development of tools for analyzing signals defined on graphs, giving rise to the field of graph signal processing (GSP) [1, 2, 3, 4, 5, 6].
@@
GSP relies on graph topology, extending traditional discrete signal processing to signals characterized by the complex and irregular underlying structures represented by corresponding graphs. Emphasizing the relationship, interaction, and effects between signals and their graph structures in various real-world scenarios, GSP aims to build a theoretical framework for high-dimensional and large-scale data analysis tasks from a signal processing perspective. This includes graph transforms [1, 2, 3, 4], frequency analysis [7, 8], filtering [9, 10], sampling and interpolation [11, 12, 13, 14, 15, 16, 17], reconstruction and recovery [18, 19, 20], uncertainty principles on graphs [12, 21, 22], fast computation algorithms [23, 14], and more. In GSP, spectral analysis plays a crucial role, inspired by harmonic analysis, using the graph Laplacian operator as its core theoretical foundation, naturally extending concepts such as frequency and filter banks to the graph domain [1, 3]. On the other hand, inspired by algebraic methods, the multiplication of the graph signal with the adjacency matrix of the underlying graph yields the fundamental shift operation for graph signals [2, 4]. These two approaches bring complementary perspectives and tools, collectively fostering attention and development in the GSP field. Although the adjacency matrix does not directly represent signal variations as the Laplacian operator does, its eigenvectors still reveal the smoothness and variation characteristics of signals on the graph [7]. In this paper, we adopt an algebraic method framework to broaden its scope.
@@
A crucial tool in this framework is the graph Fourier transform (GFT), which expands signals into eigenvectors of the adjacency matrix, defining the spectrum through corresponding eigenvalues. It is a direct generalization of the discrete Fourier transform to graphs [4]. Although the GFT can perform various operations on real-world graph signals, it cannot capture the transformation process from the vertex domain to the spectral domain and cannot handle graph signals with chirp-like characteristics [24, 25]. In recent years, approaches like windowed graph Fourier transform [8], graph fractional Fourier transform (GFRFT) [24], windowed graph fractional Fourier transform [26], and the directed graph fractional Fourier transform [27] have been proposed to address issues with these methods. However, these approaches still suffer from insufficient degrees of freedom, lack of flexibility, and underutilization of parameters. In our previous study [28], we extended the linear canonical transform (LCT) [29] to graphs, introducing the graph linear canonical transform (GLCT). As the generalization of the Fourier transform and fractional Fourier transform, the GLCT, with its three free parameters, exhibits strong adaptability and flexibility in signal processing. It demonstrates clear advantages in handling non-stationary signals and non-integer-order cases. We prove that it satisfies all expected properties, unifying GFT, GFRFT, and graph scale transform, and provide two examples of GLCT.
@@
In this paper, we further explore the uncertainty principle in the GLCT domain. The Heisenberg uncertainty principle in continuous-time signals [30], highlights the fundamental trade-off between the signal’s spread in time and its spread in frequency. Specifically, continuous-time signals cannot be perfectly localized in both time and frequency domains. In [21], the uncertainty principle for graph signals was first introduced using geodesic distance, aiming to establish a connection between the signal spread at the vertices of the graph and its spectrum spread defined by GFT in the graph spectral domain. However, a key difference between graph signals and time signals is that, while time or frequency has a well-defined distance concept, graphs are not metric spaces. Additionally, the vertices of a graph can represent more complex signals, such as those in Hilbert spaces [31]. To overcome the complexity of defining distances, M. Tsitsvero, and others [12] creatively introduced prolate spheroidal wave functions [32, 33] into graph signals, defining a new graph uncertainty principle using the energy percentage. We leverage this approach to derive the uncertainty principle in the GLCT domain. Leveraging this approach, we delineate the uncertainty principle in the GLCT domain, which proves to be broader in scope compared to those in the GFT domain. As the parameters of the GLCT undergo modifications, the scope of the uncertainty principle correspondingly shifts. This dynamic interplay furnishes pivotal insights for the formulation of the sampling theorem in the GLCT domain.
Furthermore, based on the proposed uncertainty principles of GLCT and its localization properties, we consider sampling and reconstruction in the GLCT domain. In [11], it was demonstrated that bandlimited graph signals can be perfectly recovered with a sufficient number of samples. There is an intriguing connection between sampling theory and uncertainty principles in the GLCT domain, leading to conditions for recovering bandlimited signals from subsets of their values. These conditions are found to be equivalent to those derived in [11, 12, 13, 34]. We present several signal recovery algorithms and sampling strategies aimed at finding the optimal sampling set. Building on the relationship between the GFT, the GFRFT and, the GLCT, this paper establishes a framework for graph signal sampling based on prior information in the GLCT domain. The sampling process can be viewed as a filtering operation in the GLCT domain, and the choice of filters allows the reconstruction of graph signals. In practice, our introduced sampling approaches in the GLCT domain not only address the limitations of the GFT in handling signals with chirp-like features but also demonstrates robustness in processing a variety of real-world signals. By judicious selection of optimal GLCT parameters, we can enhance performance in signal reconstruction, filtering, classification, and clustering tasks.
@@
This paper is structured as follows. Section 2 provides a brief overview of our prior work in [28], the graph uncertainty principle, and the graph sampling framework. In Section 3, based on conditions for perfect localization, we propose uncertainty principles in the GLCT domain and discuss their relationship with sampling. In Section 4, we begin by defining $\mathbf{M}$-bandlimited signals, introduce conditions for sampling and perfect recovery in the GLCT domain, and select the optimal sampling operator from qualified sampling operators. We present various sampling strategies and compare them. Section 5 evaluates the classification performance of GLCT sampling on online blogs and clustering of IEEE 118 bus test cases, comparing it with GFT and GFRFT sampling. Finally, Section 6 concludes this paper.
@@ 2 Preliminaries
In this section, we briefly review concepts in discrete signal processing on GLCT [28] and introduce the uncertainty principle of graph signals [21] and the basic theory of sampling [11, 12, 13].
2.1 Graph Linear Canonical Transform
In discrete signal processing on graphs, the high-dimensional structure of a signal is represented by a graph $\mathcal{G}=(\mathcal{V},\mathbf{A}),$ where $\mathcal{V}=\{v_{0},\dots,v_{N-1}\}$ denotes the set of vertices and $\mathbf{A}\in\mathbb{C}^{N\times N}$ denotes the weighted adjacency matrix. The graph signal is defined as a mapping that maps the vertex $v_{n}$ to a signal coefficient $x_{n}\in\mathbb{C}$, which can also be written as a complex vector
$\mathbf{x}=\begin{bmatrix}x_{0}&x_{1}&\ldots&x_{N-1}\end{bmatrix}^{\top}\in\mathbb{C}^{N}.$
The GFT can be defined through the eigendecomposition of either the adjacency matrix or the Laplacian matrix [1, 2, 3, 4]. Given that the discrete Fourier transform is equivalent to the GFT in the context of cyclic graphs (directed graphs) [2, 4], we opt for the adjacency matrix to broaden the scope of application and derive the equivalent transformation of the discrete linear canonical transform (DLCT) [35] on graphs. Hence, the adjacency matrix
$\mathbf{A=V\Lambda_{A}V}^{-1}=\sum^{N-1}_{i=0}\lambda_{\mathbf{A}_{i}}u_{i}u^{\ast}_{i},$ (1)
where the columns $u_{0},u_{1},\dots,u_{N-1}$ of $\mathbf{V}$ are the eigenvectors of $\mathbf{A}$, and the eigenvalues in the corresponding diagonal matrix $\mathbf{\Lambda_{A}}\in\mathbb{C}^{N\times N}$ are $\lambda_{\mathbf{A}_{0}},\dots,\lambda_{\mathbf{A}_{N-1}}$. Consequently, the GFT is formulated as $\mathbf{\hat{x}}=\mathbf{V}^{-1}\mathbf{x}$.
Analogous to how the GFT extends the DFT, we propose the extension of the DLCT [35] to GSP, introducing the GLCT. The definition of GLCT can be divided into three stages: the GFRFT, graph scaling, and graph chirp multiplication (CM) [28].
  1. 1.
    GFRFT Stage: By performing eigendecomposition on the orthogonal and diagonalized GFT matrix $\mathbf{V}^{-1}$, the GFRFT operator can be obtained as [24]
    $\mathbf{F^{\alpha}}=\mathbf{Q}\mathbf{\Lambda}^{\alpha}\mathbf{Q}^{\top},$ (2)
@@ 2.1 Graph Linear Canonical Transform
  2. 2.
    Graph Scaling Stage: We consider the scaled adjacency matrix $\mathbf{S}=\frac{1}{\beta}\mathbf{A}$ as the graph shift operator [9]. Similar to the GFT, by performing eigendecomposition on $\mathbf{S}$, we obtain the scaled GFT matrix $\mathbf{V}^{-1}_{\beta}$, from which further eigendecomposition yields
@@
  3. 3.
    Graph CM Stage: Similar to the definition of the discrete chirp-Fourier transform [36], the CM is performed on the GFT matrix $\mathbf{V}^{-1}$, as
@@
    where $\xi=lk+f$ is the CM parameter which is a linear function of $k$, with $l$ and $f$ being constants, and the matrix $\mathbf{\Lambda}^{\xi}$ is $\mathrm{diag}(\lambda^{\xi}_{k})=\mathrm{diag}(\exp\left(j\frac{\pi}{2}k\xi\right))$.
Based on the three stages described above, according to Eqs (2)-(4), we can derive the eigendecomposition of the GLCT as depicted in Figure 1. Its definition is as follows.
Figure 1: Eigendecomposition of the GLCT [28].
@@
$\xi=\frac{ac+bd}{a^{2}+b^{2}},\ \ \beta=\sqrt{a^{2}+b^{2}},\ \ \alpha=\cos^{-1}\left(\frac{a}{\beta}\right)=\sin^{-1}\left(\frac{b}{\beta}\right).$ (6)
@@ 2.2 Graph Uncertainty Principle
The above uncertainty principle studies the trade-off between the graph signal distribution and its spectral domain based on the specific definition of graph distances. However, a key distinction between the graph signal and the time signal is that time exists in a metric space with a well-defined distance concept, whereas the graph is not a metric space, and, its vertices can be assigned mathematical objects endowed with a richer structure, rather than assigning real numbers to each vertex. An example of such mathematical objects originates from Hilbert spaces [31]. In this case, defining distances between vertices is not straightforward.
@@
This method characterizes temporal spread based on the percentage of energy falling within the interval $[-T/2,T/2]$, defined as
$\frac{\int^{T/2}_{-T/2}\left|x\left(t\right)\right|^{2}\mathrm{d}t}{\int^{+\infty}_{-\infty}\left|x\left(t\right)\right|^{2}\mathrm{d}t}=\zeta^{2}.$ (9)
Similarly, the definition of frequency domain spread is given by
$\frac{\int^{W/2}_{-W/2}\left|\hat{x}\left(w\right)\right|^{2}\mathrm{d}w}{\int^{+\infty}_{-\infty}\left|\hat{x}\left(w\right)\right|^{2}\mathrm{d}w}=\eta^{2}.$ (10)

### v3

@@ 1 Introduction
Advancements in information and communication technologies have led to the integration of various types of data in fields such as social networks, recommendation systems, medical image processing, and bioinformatics. Unlike traditional time series or digital images, these datasets often exhibit complex and irregular structures, posing challenges for conventional signal processing tools [1, 2]. Over the past decade, significant progress has been made in the development of tools for analyzing signals defined on graphs, giving rise to the field of graph signal processing (GSP) [1, 2, 3, 4, 5].
@@
GSP relies on graph topology, extending traditional discrete signal processing to signals characterized by the complex and irregular underlying structures represented by corresponding graphs. Emphasizing the relationship, interaction, and effects between signals and their graph structures in various real-world scenarios, GSP aims to build a theoretical framework for high-dimensional and large-scale data analysis tasks from a signal processing perspective. This includes graph transforms [1, 2, 3], frequency analysis [6, 7], filtering [8, 9], uncertainty principles on graphs [12, 10, 11], sampling and interpolation [12, 13, 14, 15, 16, 17, 18, 19, 20], reconstruction and recovery [22, 23, 24, 25, 21, 26], fast computation algorithms [15, 27, 28], and more. In GSP, spectral analysis plays a crucial role, inspired by harmonic analysis, using the graph Laplacian operator as its core theoretical foundation, naturally extending concepts such as frequency and filter banks to the graph domain [1]. On the other hand, inspired by algebraic methods, the multiplication of the graph signal with the adjacency matrix of the underlying graph yields the fundamental shift operation for graph signals [2, 3]. These two approaches bring complementary perspectives and tools, collectively fostering attention and development in the GSP field. Although the adjacency matrix does not directly represent signal variations as the Laplacian operator does, its eigenvectors still reveal the smoothness and variation characteristics of signals on the graph [6]. In this paper, we adopt an algebraic method framework to broaden its scope.
@@
A crucial tool in this framework is the graph Fourier transform (GFT), which expands signals into eigenvectors of the adjacency matrix, defining the spectrum through corresponding eigenvalues. It is a direct generalization of the discrete Fourier transform (DFT) to graphs [3]. Although the GFT can perform various operations on real-world graph signals, it cannot capture the transformation process from the vertex domain to the spectral domain and cannot handle graph signals with chirp-like characteristics [29, 30]. In recent years, approaches like windowed graph Fourier transform [7], graph fractional Fourier transform (GFRFT) [29], windowed graph fractional Fourier transform [31], and the directed graph fractional Fourier transform [32] have been proposed to address issues with these methods. However, these approaches still suffer from insufficient degrees of freedom, lack of flexibility, underutilization of parameters, and inability to effectively handle non-stationary signals and non-integer-order situations.
Inspired by the handling of nonlinear and non-stationary signals in discrete signal processing (DSP), we also introduce the discrete linear canonical transform (DLCT) [33] that can solve the limitations of the DFT [34] and discrete fractional Fourier transform (DFRFT) [35] into the GSP. Generalizing the DFT to the DLCT is not a straightforward process; it presents significant challenges. This generalization encompasses a range of transformations including the DFT, DFRFT, Scaling, Fresnel Transform, and chirp multiplication (CM). Consider a chirp signal characterized by a frequency variation over time, as Fig. 1, which is prevalent in various engineering and scientific domains, including radar, communications, and medical imaging applications [33, 34, 35, 36]. Since DFT treats the entire signal as a stationary signal and cannot capture temporal dynamics, its magnitude spectrum lacks detailed information about how the frequency changes over time. In contrast, DLCT shows excellent adaptability to such nonlinear and non-stationary signals, and can more clearly distinguish the frequency changes of the signal while having higher frequency resolution.
Figure 1: The amplitude of signal $x(t)=\exp(2j\pi(f_{0}t+0.5\mu t^{2}))$ and its spectrum after DFT and DLCT of $[0.1483,-0.9889;0.9889,0.1483]$, with sampling frequency $f_{s}=1000$ Hz, time range $t=[0,1]$, initial frequency $f_{0}=50$ Hz, and frequency change rate $\mu=150$.
In our previous research [37], we introduced the graph linear canonical transform (GLCT) as a generalization of the DLCT in GSP. Serving as an extension of the GFT and the GFRFT, the GLCT, characterized by its three free parameters, exhibits strong adaptability and flexibility in signal processing. It demonstrates clear advantages in handling non-stationary signals and non-integer-order cases. We prove that it satisfies all expected properties, unifying GFT, GFRFT, graph chirp transform, and graph scale transform, and provide two examples of GLCT.
@@
In this paper, we further explore the uncertainty principle in the GLCT domain. The Heisenberg uncertainty principle in continuous-time signals [38], highlights the fundamental trade-off between the signal’s spread in time and its spread in frequency. In [10], the uncertainty principle for graph signals was first introduced using geodesic distance, aiming to establish a connection between the signal spread at the vertices of the graph and its spectrum spread defined by GFT in the graph spectral domain. However, a key difference between graph signals and time signals is that, while time or frequency has a well-defined distance concept, graphs are not metric spaces. Additionally, the vertices of a graph can represent more complex signals, such as those in Hilbert spaces [39]. To overcome the complexity of defining distances, M. Tsitsvero et al. [12] creatively introduced prolate spheroidal wave functions [40, 41] into graph signals, defining a new graph uncertainty principle using the energy percentage. Leveraging this approach, we delineate the uncertainty principle in the GLCT domain, which proves to be broader in scope compared to those in the GFT domain. As the parameters of the GLCT undergo modifications, the scope of the uncertainty principle correspondingly shifts. This dynamic interplay furnishes pivotal insights for the formulation of the sampling theorem in the GLCT domain.
Furthermore, based on the proposed uncertainty principles of GLCT and its localization properties, we consider sampling and reconstruction in the GLCT domain. Most existing methods focus on smooth or bandlimited signals, where the signal’s energy tends to be concentrated in a subset of the eigenvectors of the Laplacian or adjacency matrix [12, 13, 14, 15, 16, 17, 18, 19]. Therefore, sampling becomes a problem of selecting the optimal rows of the eigenvector matrix corresponding to the GFT [6, 12, 13, 14, 15, 16, 17]. Investigations into the relationship between signal spread on graphs and its spectral spread [12] have been proposed, along with optimal sampling techniques for signals with known frequency support [13]. Methods for noise-robust recovery and blue noise sampling have been developed [21]. In addition, a novel scalable sampling reconstruction method with parallelization has been proposed [22], along with non-Bayesian or variational Bayesian estimators based on the Cramer-Rao bound for signal recovery [23, 24]. Joint sampling and reconstruction of time-varying signals on directed graphs and the graph signal generalized sampling with prior information based on the GFRFT are proposed in [25] and [18, 19] respectively, with most of these bandlimited signal sampling methods employing greedy algorithms to select the optimal sampling sets. There is an intriguing connection between sampling theory and uncertainty principle in the GLCT domain, leading to conditions for recovering bandlimited signals from subsets of their values. We present several signal recovery algorithms and sampling strategies aimed at finding the optimal sampling set. Building on the relationship between the GFT, the GFRFT and, the GLCT, this paper establishes a framework for graph signal sampling based on prior information in the GLCT domain.
@@
This paper is structured as follows. Section 2 provides a brief overview of our prior work in [37], the graph uncertainty principle, and the graph sampling framework. In Section 3, based on conditions for perfect localization, we propose the uncertainty principle in the GLCT domain and discuss their relationship with sampling. In Section 4, we begin by defining $\mathbf{M}$-bandlimited signals, introduce conditions for sampling and perfect recovery in the GLCT domain, and select the optimal sampling operator from qualified sampling operators. We present various sampling strategies and compare them. Section 5 evaluates the classification performance of GLCT sampling on online blogs and clustering of IEEE 118 bus test cases, comparing it with GFT and GFRFT sampling. Finally, Section 6 concludes this paper. [^1] Code Available in: https://github.com/Zhangyubit/GLCTsampling
@@ 2 Preliminaries
In this section, we briefly review concepts in DSP on the GLCT [37] and introduce the uncertainty principle of graph signals [10] and the basic theory of sampling [12, 13, 14, 15, 16].
2.1 Discrete Linear Canonical Transform on Graphs
In DSP on graphs, the high-dimensional structure of a signal is represented by a graph $\mathcal{G}=(\mathcal{V},\mathbf{A}),$ where $\mathcal{V}=\{v_{0},\dots,v_{N-1}\}$ denotes the set of vertices and $\mathbf{A}\in\mathbb{C}^{N\times N}$ denotes the weighted adjacency matrix. The graph signal is defined as a mapping that maps the vertex $v_{n}$ to a signal coefficient $x_{n}\in\mathbb{C}$, which can also be written as a complex vector $\mathbf{x}=\left[x_{0}\ x_{1}\ \ldots\ x_{N-1}\right]^{\top}\in\mathbb{C}^{N}$.
1) Graph Fourier Transform: It can be defined through the eigendecomposition of either the adjacency matrix or the Laplacian matrix [1, 2, 3]. Given that the DFT is equivalent to the GFT in the context of cyclic graphs (directed graphs) [2, 3], we opt for the adjacency matrix to broaden the scope of application and derive the equivalent transformation of the DLCT [33] on graphs. Consequently, the GFT is defined as $\mathbf{\hat{x}}=\mathbf{V}^{-1}\mathbf{x}$, where $\mathbf{V}^{-1}$ is the inverse of the matrix composed of the eigenvectors obtained from the eigendecomposition of the adjacency matrix $\mathbf{A}$. Analogous to how the GFT extends the DFT, we propose the extension of the DLCT to GSP, introducing the GLCT. Utilizing the center discrete dilated Hermite functions eigendecomposition [33], the DLCT can be decomposed into three stages: the DFRFT, scaling, and CM. Therefore, the definition of the GLCT also encompasses three stages: the GFRFT, graph scaling, and graph CM [37].
2) Graph Fractional Fourier Transform: By performing eigendecomposition on the orthogonal and diagonalized GFT matrix $\mathbf{V}^{-1}$, the GFRFT operator can be obtained as [29, 30]
$\mathbf{F^{\alpha}}=(\mathbf{V}^{-1})^{\alpha}=\mathbf{Q}\mathbf{\Lambda}^{\alpha}\mathbf{Q}^{\top},$ (1)
@@ 2.1 Discrete Linear Canonical Transform on Graphs
3) Graph Scaling: Inspired by the concept of “graph time” obtained through the discrete ordered diffusion of the graph shift operator, we consider the scaled adjacency matrix $\mathbf{S}=\frac{1}{\beta}\mathbf{A}$ as the graph scaling operator [8]. Similar to the GFT, by performing eigendecomposition on $\mathbf{S}$, we obtain the scaled GFT matrix $\mathbf{V}^{-1}_{\beta}$, from which further eigendecomposition yields
@@
4) Graph Chirp Multiplication: Similar to the definition of the discrete chirp-Fourier transform [42], the graph CM is performed on the GFT matrix $\mathbf{V}^{-1}$, as
@@
where $\xi=lk+f$ is the CM parameter which is a linear function of $k$, with $l$ and $f$ being constants, and the matrix $\mathbf{\Lambda}^{\xi}$ is $\mathrm{diag}(\lambda^{\xi}_{k})=\mathrm{diag}(\exp\left(j\frac{\pi}{2}k\xi\right))$. Therefore, extending the GFT to the GLCT poses certain challenges.
Remark 1.
Theoretically, in traditional DSP, the DLCT typically comprises at least three components: the DFRFT, scaling transform, and CM. Consequently, in GSP, the GLCT cannot be as straightforwardly extended from the GFT as the GFRFT. Instead, it necessitates the proposal and identification of new graph-based counterparts for scaling transform and CM.
5) Graph Linear Canonical Transform: Based on the aforementioned four transforms and Eqs. (1)-(3), we summarize their interrelationships in Fig. 2 and define the GLCT as follows.
Figure 2: Relationship between GLCT and other related transforms.
@@
$\xi=\frac{ac+bd}{a^{2}+b^{2}},\ \ \beta=\sqrt{a^{2}+b^{2}},\ \ \alpha=\frac{2}{\pi}\cos^{-1}\left(\frac{a}{\beta}\right)=\frac{2}{\pi}\sin^{-1}\left(\frac{b}{\beta}\right).$ (5)
@@
An example is given using the Minnesota road graph [37], as shown in Fig. 3. Panel (a) depicts the original signal $\mathbf{x}=\mathbf{O^{M}}\cdot[1,1,...,-1,-1]^{\top}$ is a sequence of length $N$, panel (b) and (c) show the spectrum after the GFT and the GLCT, respectively. By adjusting the parameters of the GLCT, different perspectives of the graph spectrum can be obtained, similar to the DLCT, which can make the spectrum more concentrated.
(a) Original graph signal.
(b) Spectrum after the GFT.
(c) Spectrum after the GLCT.
Figure 3: GLCT of bipolar rectangular signals with different parameter matrices on the Minnesota road map, where the parameter matrix of (a) is $[1,0;0,1]$, (b) is $[0,1;-1,0]$, and (c) is $[\sqrt{2}/2,\sqrt{2}/2;-\sqrt{2}/4,3\sqrt{2}/4]$.
@@ 2.2 Graph Uncertainty Principle
The above uncertainty principle studies the trade-off between the graph signal distribution and its spectral domain based on the specific definition of graph distances. However, a key distinction between the graph signal and the time signal is that time exists in a metric space with a well-defined distance concept, whereas the graph is not a metric space, and, its vertices can be assigned mathematical objects endowed with a richer structure, rather than assigning real numbers to each vertex. An example of such mathematical objects originates from Hilbert space [39]. In this case, defining distances between vertices is not straightforward.
@@
The method characterizes temporal and frequency domain spread based on the percentage of energy falling within the interval $[-T/2,T/2]$ and $[-W/2,W/2]$, respectively, defined as
$\frac{\int^{T/2}_{-T/2}\left|x\left(t\right)\right|^{2}\mathrm{d}t}{\int^{+\infty}_{-\infty}\left|x\left(t\right)\right|^{2}\mathrm{d}t}=\zeta^{2},\text{and}\ \ \frac{\int^{W/2}_{-W/2}\left|\hat{x}\left(w\right)\right|^{2}\mathrm{d}w}{\int^{+\infty}_{-\infty}\left|\hat{x}\left(w\right)\right|^{2}\mathrm{d}w}=\eta^{2}.$ (8)